Two small conducting and identical spheres A and B have charges -32nC and +20nC respectively, they are separated by a distance of 0.23m.


a. What is the magnitude of the electric force between two spheres? Is this force attractive or repulsive?




b. What is the magnitude of the gravitational force between two spheres? Is this force attractive or repulsive?


1
Expert's answer
2022-03-13T18:46:09-0400

a. The magnitude of the electric force is:



"F = k\dfrac{|q_1q_2|}{r^2}\\
F = 9\times 10^9\cdot \dfrac{32\times 10^{-9}\cdot 20\times 10^{-9}}{0.23^2} \approx 1.1\times 10^{-4}N"

The force is attractive, since the charges have opposite signs.


b. The magnitude of the gravitational force can not be found, since it depends on the masses of the spheres. Gravitaional force is always attractive.